Monark Corp. just paid an annual dividend of $1.50 and expects to pay dividend of $1.80 next year, $2.20 in two years, and $2.55 in three years. Subsequently, Monark expects dividends to grow at a rate consistent with its expected earnings retention/investment rate of 60% and its expected realized return on equity of 15%. The market requires a return of 12% on Monark stock. What is your best estimate of Monark's current stock price and stock price three years from now?
Answer #1

Growth =ROE*Retention rate =15%*60% =9%
Stock three years from now =Dividend Year 3*(1+growth)/(Required Rate-Growth) =2.55*(1+9%)/(12%-9%) =92.65

Stock Current Price now=D1/(1+rate)+D2/(1+Rate)^2+Dividend Year 3/(1+Rate)^3+Stock Price 3/(1+Rate)^3
=1.80/(1+12%)+2.20*(1+12%)^2+2.55/(1+12%)^3+92.65/(1+12%)^3 =71.12